# RASPNet: A Benchmark Dataset for Radar Adaptive Signal Processing Applications

### v2

@@ 5 Limitations and Other Applications
Apart from target localization, RASPNet serves as a comprehensive testbed for benchmarking adaptive radar target detection and classification algorithms. While it currently does not support target tracking capabilities, we plan to incorporate this feature in future work. A central limitation of RASPNet is the substantial size of its clutter datasets, with the entire dataset exceeding 16 TB and each scenario’s clutter data surpassing 160 GB. These large sizes are necessary to ensure sufficient resolution in range and Doppler and to provide ample clutter realizations for data-driven methods.
@@ 6 Conclusion
In this work, we presented RASPNet, a large-scale dataset for radar adaptive signal processing applications, consisting of $100$ scenarios across the contiguous United States with $10{,}000$ clutter realizations each. RASPNet provides a standardized benchmark for evaluating and developing RASP algorithms and complex-valued learning algorithms, bridging a significant gap in the availability of a comprehensive, realistic dataset for adaptive radar processing applications. We outlined relevant properties of RASPNet, described its organization, and provided two applications of the dataset for benchmarking target localization and transfer learning in realistic scenarios. Our experiments demonstrated RASPNet’s capability to support both traditional and data-driven approaches, highlighting its potential to accelerate research and development in the adaptive radar community.
Acknowledgements
@@ Appendix B MCARM Radar Platform Parameters
The list of parameters pertaining to the airborne radar platform from the MCARM measured dataset are provided in this section (Babu et al., 1996). To simulate the clutter realizations for the RFView and bald-earth cases, we instantiate an airborne radar platform in RFView® utilizing the parameters provided in Table 2. The complete description of MCARM parameters is provided in (Himed, 1999).
@@
[…]
@@
[…]
Number of receiver channels $(\tilde{L})$horizontal $\times$ vertical} $24\times 1$
[…]
Number of pulses $(\tilde{\Lambda})$ $128$
@@
Total number of range bins $(\tilde{G})$ $630$
@@ C.1 Radar Processing Region Parameters
Sub-region extrema range bin midpoints $(r_{\text{min}},r_{\text{max}})$ $(10851\ \text{m},11451\ \text{m})$
@@
Sub-region extrema range bin midpoints $(r_{\text{min}},r_{\text{max}})$ $(11073\ \text{m},11673\ \text{m})$
@@
Sub-region extrema range bin midpoints $(r_{\text{min}},r_{\text{max}})$ $(11471\ \text{m},12071\ \text{m})$
@@
Sub-region extrema range bin midpoints $(r_{\text{min}},r_{\text{max}})$ $(11388\ \text{m},11988\ \text{m})$
@@
Sub-region extrema range bin midpoints $(r_{\text{min}},r_{\text{max}})$ $(11381\ \text{m},11981\ \text{m})$
@@ D.1 Regression CNN
The input data first passes through a convolutional layer with a $3\times 1$ kernel and stride of 1, producing 32 feature maps, followed by ReLU activation and batch normalization. Max pooling is then applied with a $2\times 1$ kernel. This output goes through another convolutional layer with a $3\times 1$ kernel and stride of 1, producing 64 feature maps, again followed by ReLU activation, batch normalization, and max pooling utilizing a $2\times 1$ kernel. The output is flattened and passed through two fully connected layers to estimate the target location $\smash{(\hat{r},\hat{\theta})}$ in range-azimuth coordinates, which is then converted to Cartesian coordinates $\smash{(\hat{x},\hat{y})}$.
D.2 Steinmetz Neural Network
The Steinmetz neural network is designed to handle both real and imaginary components of the input data separately before combining them for the final prediction. This architecture is evaluated on the CVNN benchmark from Section 4.2.1. The analytic neural network is a Steinmetz neural network in which the separately processed real and imaginary components are related by the discrete Hilbert transform (DHT), weighted by a tradeoff parameter, $\beta$. Across all experiments in Section 4.2.1, we let $\beta=1e{-}3$, and consider $lN=128$, $k=2$ (see below).
@@ D.2 Steinmetz Neural Network
D.3 Real-Valued Neural Network
The real-valued neural network (RVNN) architecture is an approach for handling both real and imaginary components by concatenating them and processing them together. This architecture is evaluated on the CVNN benchmark from Section 4.2.1. Across all experiments in Section 4.2.1, we consider $lN=128$, $k=2$ (see below).
@@ D.3 Real-Valued Neural Network
D.4 Complex-Valued Neural Network
The Complex-Valued Neural Network (CVNN) is designed to handle complex-valued data by treating the real and imaginary parts jointly as complex numbers. To implement this network, we use the Complex PyTorch library (Matthès et al., 2021). This architecture is evaluated on the CVNN benchmark from Section 4.2.1. Across all experiments in Section 4.2.1, we consider $lN=128$, $k=2$ (see below).
@@ E.1 Dataset Access and Organization
RASPNet is hosted within the Duke University Compute Cluster and can be accessed using the following Globus link: https://duke.is/8/jvkg. The dataset consists of $100$ radar scenarios from across the contiguous United States, each comprising $10{,}000$ clutter realizations. The main webpage for the dataset can be accessed at: https://shyamven.github.io/RASPNet.
The base folder in RASPNet contains three folders: EXAMPLES, CVNN, and CLUTTER. The EXAMPLES folder consists of data files used to generate the non-CVNN results presented in the main text, the CVNN folder consists of data files used to generate the CVNN results presented in the main text, and the CLUTTER folder consists of $100$ subfolders, each of which pertain to a different radar scenario. These scenarios are ordered from Westernmost to Easternmost, such that the folder CLUTTER/num1_lat_lon/ is the Westernmost scenario in RASPNet, and the folder CLUTTER/num100_lat_lon/ is the Easternmost scenario in RASPNet (lat and lon denote the platform latitude and longitude for each scenario). This nomenclature follows from the main text, where $M=100,S=\{1,2,\ldots,M\}$, and $i\in S$.
[…]
@@ E.2 Code Access
The code files for generating the empirical results provided within the main text can be found in the attached supplementary .zip file. We organize these files into two separate subfolders: Localization_Examples and Transfer_Learning_Examples. The Localization_Examples consist of code files for benchmarking non-CVNN and CVNN target localization accuracy in the scenarios $i\in\{29,35,60,62,76\}$. The Transfer_Learning_Examples consist of example transfer learning code files, where we train a regression neural network on scenario $i=29$, and evaluate the network for $i\in\{35,60,62,76\}$.
@@ F.1 Motivation
    Who funded the creation of the dataset? The main funding body is the Air Force Office of Scientific Research.
@@ F.2 Distribution
    How will the dataset will be distributed (e.g., tarball on website, API, GitHub)? The dataset can be downloaded using the provided Globus link and is hosted on the Duke University Compute Cluster.
@@ F.3 Maintenance
    Who will be supporting/hosting/maintaining the dataset? The authors, with the Duke University Office of Information Technology, will support, host, and maintain the dataset.
@@
    Is there an erratum? No. If errors are found in the future, we will release errata on the main webpage for the dataset.
@@
    Will the dataset be updated (e.g., to correct labeling errors, add new instances, delete instances)? Yes, the datasets will be updated whenever necessary to ensure accuracy, and announcements will be made accordingly. These updates will be posted on the main webpage for the dataset.
@@
    Do the authors bear all responsibility in case of violation of rights, and confirm the data license?) Yes, the authors bear all responsibility in case of violation of rights.
@@ F.4 Composition
    What do the instance that comprise the dataset represent (e.g., documents, photos, people, countries?) RASPNet comprises $100$ airborne radar scenarios from across the contiguous United States. Each scenario consists of $10{,}000$ clutter realizations, which are stored as multidimensional matrices (.mat files).
@@
    How many instances are there in total (of each type, if appropriate)? There are $1{,}000{,}000$ total clutter realizations present within RASPNet, amounting to more than $16$ TB of data.
@@
  10. 10.
    Does the dataset contain data that, if viewed directly, might be offensive, insulting, threatening, or might otherwise cause anxiety? No.
@@ G.1 Radar Platform Parameters
[…]
@@ G.2 Basic Scenarios
[Uncaptioned image] Platform Aim Location (lat, lon): $(37.8459^{\circ},-97.7250^{\circ})$ Platform Position (lat, lon): $(37.9689^{\circ},-97.5750^{\circ})$ Folder: /CLUTTER/num60_lat37.9689_lon-97.575/ Description: Scenario $q_{i}=4\ (i=60)$ in RASPNet is in the Great Plains region of Central Kansas. The radar platform is positioned at latitude $37.9689^{\circ}$ N and longitude $-97.5750^{\circ}$ W. It aims towards a location at latitude $37.8459^{\circ}$ N and longitude $-97.7250^{\circ}$ W. This region is characterized by an expansive and flat grassland environment, with low terrain variability.
@@ G.3 Intermediate Scenarios
[Uncaptioned image] Platform Aim Location (lat, lon): $(41.9734^{\circ},-80.3229^{\circ})$ Platform Position (lat, lon): $(42.0964^{\circ},-80.1729^{\circ})$ Folder: /CLUTTER/num84_lat42.0964_lon-80.1729/ Description: Scenario $q_{i}=28\ (i=84)$ in RASPNet is on the Lake Erie coastline of Pennsylvania. The radar platform is positioned at latitude $42.0964^{\circ}$ N and longitude $80.1729^{\circ}$ W. It aims towards a location at latitude $41.9734^{\circ}$ N and longitude $80.3229^{\circ}$ W. The radar is positioned near the coastline, the aim direction is inland, with nonuniform terrain.
@@
Scenario $q_{i}=48$: Sandia Crest, NM ($i=48)$
[Uncaptioned image] Platform Aim Location (lat, lon): $(38.7045^{\circ},-105.0988^{\circ})$ Platform Position (lat, lon): $(38.8275^{\circ},-104.9488^{\circ})$ Folder: /CLUTTER/num48_lat38.8275_lon-104.9488/ Description: Scenario $q_{i}=48\ (i=48)$ in RASPNet is in the Sandia Mountains of New Mexico. The radar platform is positioned at latitude $38.8275^{\circ}$ N and longitude $104.9488^{\circ}$ W. It aims towards a location at latitude $38.7045^{\circ}$ N and longitude $105.0988^{\circ}$ W. This region is noted for its variable terrain, and the radar is positioned behind a mountain crest.

### v3

@@ 5 Limitations and Other Applications
Apart from the presented applications, RASPNet serves as a comprehensive testbed for benchmarking adaptive radar target detection and target classification algorithms. While it currently does not support target tracking capabilities, we plan to incorporate this feature in future work. A central limitation of RASPNet is the substantial size of its clutter datasets, with the entire dataset exceeding 16 TB and each scenario’s clutter data surpassing 160 GB. These large sizes are necessary to ensure sufficient resolution in range and Doppler and to provide ample clutter realizations for data-driven methods.
@@ 6 Conclusion
In this work, we presented RASPNet, a large-scale dataset consisting of $100$ airborne radar scenarios from across the contiguous United States. RASPNet provides a standardized benchmark for evaluating and developing RASP algorithms and complex-valued learning algorithms, bridging a significant gap in the availability of a comprehensive, realistic dataset for adaptive radar processing applications. We outlined relevant properties of RASPNet, described its organization, and provided three applications of the dataset for benchmarking target localization and CVNNs, and transfer learning in realistic scenarios. Our experiments demonstrated RASPNet’s capability to support data-driven approaches, highlighting its potential to accelerate research and development in the adaptive radar community.
Acknowledgments and Disclosure of Funding
@@ Appendix B MCARM Radar Platform Parameters
The list of parameters pertaining to the airborne radar platform from the MCARM measured dataset are provided in this section [Babu et al., 1996]. To simulate the clutter realizations for the RFView and bald-earth cases, we instantiate an airborne radar platform in RFView® utilizing the parameters provided in Table 2. The complete description of MCARM parameters is provided in [Himed, 1999].
@@
Carrier frequency $(\tilde{f}_{c})$ $1.25\,\text{GHz}$
Bandwidth (B~)($\tilde${B}) $1.25\,\text{MHz}$
Pulse Repetition Frequency $(\tilde{f}_{PR})$ $1984\,\text{Hz}$
Radar platform altitude (h~)($\tilde${h}) $3072\,\text{m}$
@@
Antenna element configuration {horizontal $\times$ vertical} $16\times 8$
Number of receiver channels (L~)($\tilde${L}) {horizontal $\times$ vertical} $24\times 1$
Antenna element horizontal spacing $0.1092\,\text{m}$
Antenna element vertical spacing $0.1407\,\text{m}$
@@
Total number of range bins (G~)($\tilde${G}) $630$
@@ C.1 Radar Processing Region Parameters
Sub-region extrema range bin midpoints $(r_{\text{min}},r_{\text{max}})$ (10851m,11451m)(10851\ $\text${m},11451\ $\text${m})
@@
Sub-region extrema range bin midpoints $(r_{\text{min}},r_{\text{max}})$ (11073m,11673m)(11073\ $\text${m},11673\ $\text${m})
@@
Sub-region extrema range bin midpoints $(r_{\text{min}},r_{\text{max}})$ (11471m,12071m)(11471\ $\text${m},12071\ $\text${m})
@@
Sub-region extrema range bin midpoints $(r_{\text{min}},r_{\text{max}})$ (11388m,11988m)(11388\ $\text${m},11988\ $\text${m})
@@
Sub-region extrema range bin midpoints $(r_{\text{min}},r_{\text{max}})$ (11381m,11981m)(11381\ $\text${m},11981\ $\text${m})
@@ D.1 Regression CNN
The input data first passes through a convolutional layer with a $3\times 1$ kernel and stride of 1, producing 32 feature maps, followed by ReLU activation and batch normalization. Max pooling is then applied with a $2\times 1$ kernel. This output goes through another convolutional layer with a $3\times 1$ kernel and stride of 1, producing 64 feature maps, again followed by ReLU activation, batch normalization, and max pooling utilizing a $2\times 1$ kernel. The output is flattened and passed through two fully connected layers to estimate the target location $\smash{(\hat{r},\hat{\theta})}$ in range-azimuth coordinates, which is then converted to Cartesian coordinates $\smash{(\hat{x},\hat{y})}$. The regression CNN has 14,91014{,}910 parameters.
D.2 Adaptive Radar Transformer
[…]
  • Conv1d Block 1:
    • conv1: 1D convolution from 21 → 32 channels, kernel size 3, padding 1.
    • bn1: BatchNorm1d over 32 channels.
    • relu1: ReLU activation.
    • pool1: MaxPool1d with kernel=2, stride=2 → length halved.
  • Conv1d Block 2:
    • conv2: 1D convolution from 32 → 64 channels, kernel size 3, padding 1.
    • bn2: BatchNorm1d over 64 channels.
    • relu2: ReLU activation.
    • pool2: MaxPool1d with kernel=2, stride=2 → length halved again.
  • Token preparation:
    • permute: Permute to (batch,L, 64)($\text${batch},\,L,\,64) where $L=26/2^{2}$.
    • token_proj: Linear layer mapping 64 → 64 to produce $L$ patch tokens.
  • CLS token & positional embeddings:
    • cls_token: Learned $(1\times 1\times 64)$ token prepended to tokens.
    • pos_embed: Learned $(1\times(L+1)\times 64)$ positional embeddings added to sequence.
  • Transformer Encoder:
    • encoder_layer: Multi‑head self‑attention ($d_{model}=64$, 4 heads) + feed‑forward ($128$ hidden), dropout 0.1.
    • transformer: Stack 2 encoder layers, producing (batch,L+1, 64)($\text${batch},\,L+1,\,64).
  • MLP head:
    • cls_rep: Extract token 0 → (batch, 64)($\text${batch},\,64).
    • ln: LayerNorm over 64 dims.
    • linear1: 64 → 32.
    • gelu: GELU activation.
[…]
D.3 Convolutional Steinmetz Network
The convolutional Steinmetz network is designed to handle both real and imaginary components of the input data separately before combining them for the final prediction. This architecture is evaluated on the CVNN benchmark from Section 4.3. The convolutional analytic network is a Steinmetz network in which the separately processed real and imaginary components are related by the discrete Hilbert transform (DHT), weighted by a tradeoff parameter, $\beta$. Across all experiments in Section 4.3, we let $\beta=1e{-}3$, and consider $lN=128$, $k=2$ (see below). The network has 368,002368{,}002 parameters.
@@ D.3 Convolutional Steinmetz Network
D.4 Convolutional Real-Valued Neural Network
The real-valued neural network (RVNN) architecture is an approach for handling both real and imaginary components by concatenating them and processing them together. This architecture is evaluated on the CVNN benchmark from Section 4.3. Across all experiments in Section 4.3, we consider $lN=128$, $k=2$ (see below). The network has 185,634185{,}634 parameters.
@@ D.4 Convolutional Real-Valued Neural Network
D.5 Convolutional Complex-Valued Neural Network
The Complex-Valued Neural Network (CVNN) is designed to handle complex-valued data by treating the real and imaginary parts jointly as complex numbers. To implement this network, we use the Complex PyTorch library [Matthès et al., 2021]. This architecture is evaluated on the CVNN benchmark from Section 4.3. Across all experiments in Section 4.3, we consider $lN=128$, $k=2$ (see below). The network has 368,388368{,}388 parameters.
@@ D.5 Convolutional Complex-Valued Neural Network
RASPNet: Supplementary Information
Shyam Venkatasubramanian
Duke University
&Bosung Kang
Air Force Research Laboratory
&Ali Pezeshki
Colorado State University
&Muralidhar Rangaswamy
Air Force Research Laboratory
&Vahid Tarokh
Duke University
@@ E.1 Dataset Access and Organization
RASPNet is hosted by the United States Air Force and can be downloaded from the AFRL Sensor Data Management System. The dataset comprises $100$ radar scenarios from across the contiguous United States, each comprising 10,00010{,}000 clutter realizations. The main webpage for the dataset can be accessed at: https://www.sdms.afrl.af.mil/index.php?collection=raspnet.
The base folder in RASPNet consists of three folders: EXAMPLES, CVNN, and CLUTTER. The EXAMPLES folder consists of data files used to generate the target localization and transfer learning results presented in the main text, the CVNN folder consists of data files used to generate the CVNN results presented in the main text, and the CLUTTER folder consists of $100$ subfolders, each of which pertain to a different radar scenario. These scenarios are ordered from westernmost to easternmost, such that the folder CLUTTER/num1_lat_lon/ is the westernmost scenario in RASPNet, and the folder CLUTTER/num100_lat_lon/ is the easternmost scenario in RASPNet (lat and lon denote the platform latitude and longitude for each scenario). This nomenclature follows from the main text, where $M=100,S=\{1,2,\ldots,M\}$, and $i\in S$.
Each of the $100$ scenario subfolders in RASPNet contain 10,00010{,}000 i.i.d. clutter realizations, where each realization is stored within $\textbf{real}\{j\}\textbf{.mat}$. The realization $\textbf{CLUTTER/num}\{i\}\textbf{\_lat\_lon/real}\{j\}\textbf{.mat}$ parallels $z^{i}_{j}$ from the main text, wherein $K=10{,}000,D=\{1,2,\ldots,K\}$, and $j\in D$. Importing the realization $\textbf{real}\{j\}\textbf{.mat}$ yields the variable X_clut, which is a size $L\times\Lambda\times G$ matrix.
The EXAMPLES folder consists of feature-labels pairs for five scenarios: $i\in\{29,35,60,62,76\}$. The features for scenario $i$ are stored in $\textbf{num}\{i\}\textbf{\_NAMF\_DATA\_25k.zip}$ and the labels are stored in $\textbf{num}\{i\}\textbf{\_Ground\_Truth\_25k.csv}$. These feature-label pairs were used to generate the target localization and transfer learning results from the main text, using the code files in Localization_Examples and Transfer_Learning_Examples (see Section E.2).
The CVNN folder consists of feature-labels pairs from five scenarios: $i\in\{29,35,60,62,76\}$. The train and test features for scenario $i$ are stored within $\textbf{num}\{i\}\textbf{.zip}$ in the folders $\textbf{num}\{i\}/\textbf{train}$ and $\textbf{num}\{i\}/\textbf{test}$, and the labels are stored in $\textbf{num}\{i\}/\textbf{train.csv}$ and $\textbf{num}\{i\}/\textbf{test.csv}$. These feature-label pairs were used to generate the CVNN target localization results from the main text using the code file RASPNet_CVNN.ipynb in Localization_Examples (see Section E.2).
@@ E.2 Code Access
The code files for generating the empirical results provided within the main text can be found at the url: https://github.com/shyamven/RASPNet. We organize these files into two separate subfolders: Localization_Examples and Transfer_Learning_Examples. The Localization_Examples consist of code files for benchmarking target localization accuracy in the scenarios $i\in\{29,35,60,62,76\}$. The Transfer_Learning_Examples consist of example transfer learning code files, where we train a regression neural network on scenario $i=29$, and evaluate the network for $i\in\{35,60,62,76\}$.
@@ F.1 Motivation
    Who funded the creation of the dataset? The main funding body is the United States Air Force Office of Scientific Research. Other funding sources of individual authors are listed in the Acknowledgments section.
@@ F.2 Distribution
    How will the dataset will be distributed (e.g., tarball on website, API, GitHub)? The dataset is hosted by the United States Air Force, and can be downloaded from the following link: (https://www.sdms.afrl.af.mil/index.php?collection=raspnet).
@@ F.3 Maintenance
    Who will be supporting/hosting/maintaining the dataset? The authors, with the United States Air Force, will support, host, and maintain the dataset.
@@
    How can the owner/curator/manager of the dataset be contacted (e.g., email address)? The owner/curator/manager of the dataset can be contacted via the following emails: Shyam Venkatasubramanian (sv222@duke.edu), and Vahid Tarokh (vahid.tarokh@duke.edu).
@@
    Is there an erratum? No. If errors are found in the future, we will release errata on the main webpage for the dataset: https://www.sdms.afrl.af.mil/index.php?collection=raspnet.
@@
    Will the dataset be updated (e.g., to correct labeling errors, add new instances, delete instances)? Yes, the datasets will be updated whenever necessary to ensure accuracy, and announcements will be made accordingly. These updates will be posted on the main webpage for the dataset: https://www.sdms.afrl.af.mil/index.php?collection=raspnet.
@@
    Do the authors bear all responsibility in case of violation of rights, and confirm the data license?) Yes, the authors bear all responsibility in case of violation of rights. The dataset, RASPNet © 2024 by Shyam Venkatasubramanian, Bosung Kang, Ali Pezeshki, Muralidhar Rangaswamy, Vahid Tarokh is licensed under CC BY 4.0.
  6. 6.
@@ F.4 Composition
    What do the instance that comprise the dataset represent (e.g., documents, photos, people, countries?) RASPNet comprises $100$ airborne radar scenarios from across the contiguous United States. Each scenario consists of 10,00010{,}000 clutter realizations, which are stored as multidimensional matrices (.mat files).
@@
    How many instances are there in total (of each type, if appropriate)? There are 1,000,0001{,}000{,}000 total clutter realizations present within RASPNet, amounting to more than $16$ TB of data.
@@ G.1 Radar Platform Parameters
Carrier frequency $(f_{c})$ $10\,\text{GHz}$
Bandwidth $(B)$ $5\,\text{MHz}$
Pulse Repetition Frequency $(f_{PR})$ $1100\,\text{Hz}$
Radar platform altitude $(h)$ $1000\,\text{m}$
Range resolution $(\Delta r)$ $30\,\text{m}$
Antenna element configuration {horizontal $\times$ vertical} $48\times 5$
@@ G.2 Basic Scenarios
[Uncaptioned image] Platform Aim Location (lat, lon): $(37.8459^{\circ},-97.7250^{\circ})$ Platform Position (lat, lon): $(37.9689^{\circ},-97.5750^{\circ})$ Folder: /CLUTTER/num60_lat37.9689_lon-97.575/ Description: Scenario $q_{i}=4\ (i=60)$ in RASPNet is in the Great Plains region of Central Kansas. The radar platform is positioned at latitude $37.9689^{\circ}$ N and longitude $97.5750^{\circ}$ W. It aims towards a location at latitude $37.8459^{\circ}$ N and longitude $97.7250^{\circ}$ W. This region is characterized by an expansive and flat grassland environment, with low terrain variability.
@@ G.3 Intermediate Scenarios
[Uncaptioned image] Platform Aim Location (lat, lon): $(41.9734^{\circ},-80.3229^{\circ})$ Platform Position (lat, lon): $(42.0964^{\circ},-80.1729^{\circ})$ Folder: /CLUTTER/num84_lat42.0964_lon-80.1729/ Description: Scenario $q_{i}=28\ (i=84)$ in RASPNet is on the Lake Erie coastline of Pennsylvania. The radar platform is positioned at latitude $42.0964^{\circ}$ N and longitude $80.1729^{\circ}$ W. It aims towards a location at latitude $41.9734^{\circ}$ N and longitude $80.3229^{\circ}$ W. The radar is positioned near the coastline, the aim direction is inland, and the surrounding terrain is nonuniform.
@@
Scenario $q_{i}=48$: Sandia Crest, NM ($i=46)$
[Uncaptioned image] Platform Aim Location (lat, lon): $(35.0164^{\circ},-106.5885^{\circ})$ Platform Position (lat, lon): $(35.1394^{\circ},-106.4385^{\circ})$ Folder: /CLUTTER/num46_lat35.1394_lon-106.4385/ Description: Scenario $q_{i}=48\ (i=46)$ in RASPNet is in the Sandia Mountains of New Mexico. The radar platform is positioned at latitude $35.1394^{\circ}$ N and longitude $106.4385^{\circ}$ W. It aims towards a location at latitude $35.0164^{\circ}$ N and longitude $106.5885^{\circ}$ W. This region is noted for its variable terrain, and the radar is positioned behind a mountain crest.